OpenStudy (anonymous):

Find the volume of a sphere with diameter 15 feet. A. 1767.15 ft3 B. 994.02 ft3 C. 235.62 ft3 D. 471.24 ft3

OpenStudy (anonymous):

\[V=\frac{4 \pi r^3}{3} \]\[\frac{4}{3}\pi \left(\frac{15}{2}\right)^3=\frac{1125 \pi }{2}=1767.15 \]

OpenStudy (anonymous):

omg thank youuuuuu

OpenStudy (anonymous):

You are welcome.